GlueX PID Meeting, July 9, 2010

From GlueXWiki
Jump to: navigation, search

Location, Time

Friday, July 9, 2010
2:00 pm EDT
JLab: CEBAF Center, Room F326/7
ESNet: 8542553
EVO: direct meeting link[1]
Conference Group: 800-377-8846, participant code: 77438230
Slides: talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2010-07 on the JLab CUE (you have to be a member of the "halld" group). This directory is accessible from the web at .


  1. Announcements
  2. Minutes from the last meeting
  3. Time-of-Flight
    1. Prototype progress [1]
    2. Status of the MOU
  4. Start Counter
    1. prototype progress
  5. Target
    1. Status of the MOU
    2. Status of refrigerator plan
  6. Action Items


FSU: Sasha Ostrovidov JLab: Eugene Chudakov, Mark Ito (chair), Yi Qiang, Sascha Somov, Elton Smith, Tim Whitlatch, Elliott Wolin, Beni Zihlmann

Review of minutes from last meeting

We went over the minutes from June 11 meeting.

PMT purchase for TOF prototypes

The PO went out. Paul Eugenio is waiting on delivery.

Refrigerator for the target

  • Tim still needs to do the cryo review. A cost estimated is needed as well.
  • Chris Keith is confident about the 4He capability of the proposed refrigerator, but this needs verification. Chris was planning to talk to the vendor about this.

Time-of-flight prototype

Sasha O. reported that the LED pulser system up and running, but has not yet been tried with the prototype counters.

Need for a laser calibration system for the time-of-flight

Elton reported on the use of a nitrogen laser in CLAS for calibration of their time-of-flight system. It was only used for for time-walk corrections and was useful but a lot of effort was necessary to build, install, and maintain the system. Time-walk corrections could have been measured with the data, but would have been difficult since the data does not have a explicity external start time. He pointed out the need to choose a frequency so that the scintillator itself is excited.

Elton was recently down at the University of South Carolina. Ralf Gothe there is doing similar things for the CLAS12 forward TOF. They have decided against a laser system; they will do everything with the data. With their cosmic ray set up they have obtained time-walk correction constants.

Elton pointed out that we will have two planes, and that will help doing the calibration with the data. A large dynamic range in energies needed. There was a lot of discussion about possible data-based calibration schemes. Clearly a study of a particular scheme or schemes is needed. If a workable scheme can be proposed, then we may not need a calibration system as elaborate as that used by CLAS.

Other TOF issues

  • Sasha mentioned that tests with a flash ADC would be desirable at some time.
  • The trapezoidal light guides are not large enough to cover PMT. That frustrated the first attempt to measure the response of the edges of the PMT's.
  • The new data acquisition system using the TI is written; it has not been tested yet.

Time-of-flight prototype progress

Sasha O. has been working on cosmic ray measurements using different wrapping material for the prototype counters. He showed plots of the ADC spectra for materials tried so far. Future plans include aluminized Mylar and ESR DF2000.

Elton mentioned that at USC, Ralf is using aluminized Mylar. That worked well for Ralf on a previous experiment. Elton also mentioned that CLAS uses aluminum foil to wrap their counters.


  1. This link is good through 8/20